For the function f(x)= x^2-2 defined over [1,5] , verify: L(P,f)≤U(-P,f) where P is the partition which divides [ 1,5] into four equal intervals.
1
Expert's answer
2021-07-12T07:26:14-0400

"L(P,f)= \sum^4_{i=1}m_i \Delta x_i\\
U(-P,f)=\sum^4_{i=1}M_i \Delta x_i\\
m_i=infimum[f(x):x_{i-1}≤x≤x_i]\\
M_i=inprenum[f(x):x_{i-1}≤x≤x_i]\\
f(1)=-1 \\
f(2)=2 \\
f(3)=7\\
f(4)=14 \\
f(5)=23 \\
Hence, m_1=-1, m_2=2, m_3=7, m_4=14\\
M_1=2, M_2=7, M_3=14, M_4=23\\
\Delta x_i=1\\
L(P,f)=-1+2+7+14=22\\
U(-P,f)=2+7+14+23=46\\
L(P,f)≤U(-P,f)\\"